# The resummation model in FLAMINGO: precisely predicting matter power suppression from observed halo baryon fractions

### v1

@@ 3.4 Contribution strength of halo baryon fractions and uncertainty propagation
Examining Fig. 8, we can immediately see that the signal is dominated by halo masses around $M_{\mathrm{500c}}=10^{14}\,\mathrm{M_{\odot}}$, in line with the findings of e.g. van Daalen et al. (2020), Debackere et al. (2020), Salcido et al. (2023) and van Loon & van Daalen (2024). The impact of other halo mass bins drops off quickly from there, with haloes in the mass range $M_{\mathrm{500c}}=10^{14.5-15}\,\mathrm{M_{\odot}}$ still contributing significantly on scales $k\,{<}\,{10}\,h\mathrm{\,Mpc^{-1}}$, while haloes with masses $M_{\mathrm{500c}}>10^{15}\,\mathrm{M_{\odot}}$ contribute negligibly on all scales. On the low-mass end, the large-scale contribution diminishes sharply below $M_{\mathrm{500c}}\approx 10^{13}\,\mathrm{M_{\odot}}$, while that on small scales ($k\,{\gtrsim}\,{5}\,h\mathrm{\,Mpc^{-1}}$) drops off much more slowly. The baryon fractions in haloes in the lowest two mass bins considered here, $M_{\mathrm{500c}}=10^{11-11.5}\,\mathrm{M_{\odot}}$ and $10^{10.5-11}\,\mathrm{M_{\odot}}$, contribute negligibly on all scales shown here. However, their inclusion in the model is still relevant to better measure the non-halo component.
@@ 3.5 Resummation based only on $\boldsymbol{\Delta=}\mathbf{500c}$
While increasingly accurate measurements of mean halo baryon fractions within $R_{\mathrm{500c}}$ are already observationally available for different $M_{\mathrm{500c}}$ (e.g. Akino et al., 2022) or even $M_{\mathrm{200c}}$ (e.g. Dev et al., 2024, though with some extrapolation), measuring these precisely for larger regions is more challenging. Promising techniques that derive large-scale gas fractions from measurements of the kSZ (e.g. Hadzhiyska et al., 2024) or tSZ (e.g. Dalal et al., 2025) signal, or those using fast radio bursts to probe the intergalactic medium (e.g. Lemos, 2025), may change this in the near future. In the meantime, it is worth investigating how well the suppression signal can be reproduced using only baryon fractions measured within $R_{\mathrm{500c}}$.
One of the results of Debackere et al. (2020) was that the suppression signal is sensitive to the density profile of hot gas outside $R_{\mathrm{500c}}$. In the terms of the resummation model, the results therefore depend on the retained mass fractions on scales that are larger than the baryon fractions were measured on. We confirm this finding by showing the difference between two extreme assumptions for where the removed mass goes in Fig. 9. On the left-hand side, we show the result of assuming that all mass removed from within $R_{\mathrm{500c}}$ stays within $R_{\mathrm{200m}}$, by setting $f_{\mathrm{ret},i,\mathrm{200m}}=1$ for all halo masses. Clearly, this is a poor assumption, and the suppression of power on large scales is strongly underestimated as a result. On the right-hand side, we instead assume that all mass removed from within $R_{\mathrm{500c}}$ also makes it out of $R_{\mathrm{200m}}$ (where it is again assumed to be distributed similar to matter already outside DMO haloes), which means that $f_{\mathrm{ret},i,\mathrm{200m}}=1-(1-f_{\mathrm{ret},i,\mathrm{500c}})b_{i,\mathrm{500c}}f_{\mathrm{M},i,\mathrm{500c}}/(b_{i,\mathrm{200m}}f_{\mathrm{M},i,\mathrm{200m}})$. While this assumption is closer to the results of the simulations, how well it works is significantly dependent on halo mass and the strength of feedback: when either very little or a lot of mass is removed from within $R_{\mathrm{500c}}$, this assumption overestimates the retained mass fraction within $R_{\mathrm{200m}}$, i.e. more mass is actually removed from within $R_{\mathrm{200m}}$ (or less of it is accreted) than just that ejected from $R_{\mathrm{500c}}$.
@@
The correlation between both retained mass fractions is particularly strong at high halo masses (large $f_{\mathrm{ret,500c}}$, see Fig. 3), where the retained mass of the larger region can generally be predicted to within percent precision from that of the smaller region. As we go to lower halo mass scales, however, the regions are increasingly decorrelated, and accurately predicting $f_{\mathrm{ret},i,\mathrm{200m}}$ (left) or $f_{\mathrm{ret},i,\mathrm{5xr500c}}$ (right) becomes more sensitive to the strength of feedback and its implementation (thermal or jet AGN feedback). Somewhat surprisingly, the correlation is slightly tighter for the larger region, $r<5\times R_{\mathrm{500c}}$ – likely because any dependence on the halo profile is removed, due to the larger radius being a fixed multiple of the smaller. This hints that we might obtain even better results if we replaced $R_{\mathrm{200m}}$ by a fixed multiple of $R_{\mathrm{500c}}$ in the resummation model, but we leave this for future work. We note that equation (17) also works well for predicting $f_{\mathrm{bc},\Delta}$ from $f_{\mathrm{bc,500c}}$, and translating these into retained mass fractions through equation (9) to predict $f_{\mathrm{ret},\Delta}$ results in a comparable scatter around the true values as directly using equation (17) for $f_{\mathrm{ret},\Delta}$.
@@
Figure 11: Similar to the bottom row of Fig. 7, but now instead of including large-scale baryon fractions, we extrapolate the small-scale information. Using the correlation between the retained fractions within $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$ (see Fig. 10 and equation 17), we predict $f_{\mathrm{ret},i,\mathrm{200m}}$ given $f_{\mathrm{ret},i,\mathrm{500c}}$. This only slightly reduces the accuracy of the model for $k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$, and allows the resummation model to be applied given only mean halo baryon fractions within $R_{\mathrm{500c}}$.
@@ 4 Discussion
The improved resummation model presented in this work is able to reproduce the $z=0$ matter power suppression signal for a wide range of FLAMINGO simulations, without any free parameters. Based only on their halo spherical overdensity masses and mean halo baryon fractions, for example $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$ or similar, the model yields percent-level precision up to $k\,{\approx}\,{10}\,h\mathrm{\,Mpc^{-1}}$, and even to $k\,{\gtrsim}\,{20}\,h\mathrm{\,Mpc^{-1}}$ if gas and stellar fractions are known inside $R_{\mathrm{2500c}}$ as well. Similar precision can be obtained at redshifts up to at least $z=1$ by adding a single parameter, although our results imply that this parameter may be independent of baryonic feedback and could therefore be a fixed function of redshift (see §3.6). We will explore this further in future work. Even without this parameter, the resummation model achieves $\lesssim 3\%$ precision up to at least $z=1$ for all $k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$.
@@ 4.1 The origin of the retained mass fraction relation
Central to the resummation model is the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation, which, for a given halo overdensity region and redshift, accurately returns the mass ratio between a halo in a hydrodynamic simulation or in the real Universe and its DMO equivalent, as a function of its halo baryon fraction (see Fig. 2, 3 and Fig. 4, as well as Appendix C). This relation is seemingly universal, showing no significant dependence on cosmology, feedback strength, or halo mass (at least for $M_{\mathrm{500c}}\gtrsim 10^{12}\,\mathrm{M_{\odot}}$). The existence of this relation, and its slope exceeding unity for all but the most massive haloes, implies that when mass is lost due to feedback or other baryonic processes, a predictable amount of cold dark matter mass is removed from overdense regions as well. As mentioned early in §3.1, this is expected based on purely gravitational processes: after mass is removed from the centre, the density enclosed in the ejection radius will decrease, and so overdensity radii will also decrease, which means mass between the new and original overdensity radii will be “lost”. Additionally, the halo potential will be lessened, reducing the binding energy of the particles, allowing some to stream out of the halo. Afterwards, a gravitational exchange of energy between halo particles will take place (halo relaxation), which will cause additional particles to move outward, or even become unbound, while for others the binding energy will increase. It is however not clear a priori whether this should lead to a precisely predictable amount of (relative) mass loss.
@@
To test whether this picture is accurate and how much freedom it allows, we ran a series of idealised DMO simulations using the swift simulation code (Schaller et al., 2024). Using ICICLE (Drakos et al., 2017), we generate initial conditions for a $M_{\mathrm{200m}}=10^{14}\,\mathrm{M_{\odot}}$ exponentially-truncated NFW halo with concentration $c\approx 4.5$ using 138 500138\,500 dark matter particles, such that $10^{5}$ of these reside within $R_{\mathrm{200m}}$. We note that the choice of halo mass is unimportant, as the mass can be rescaled without affecting the experiment. Similarly, the exponential truncation helps to define a stable “edge” to the halo, but does not influence the outcome. After letting the halo settle into a steady state, we instantaneously remove some fraction $f_{\mathrm{ej}}$ of its $M_{\mathrm{200m}}$ mass, crudely mimicking AGN feedback ejecting gas. We do so starting from the particle at the halo centre and moving radially outward until the desired fraction is reached, then remove these particles from the simulation. [^13] We also tried choosing some inner radius, and removing a random subset of the particles inside it, corresponding to the mass fraction to be “ejected”. The results are indistinguishable. We then continue the simulation. Immediately after this idealised ejection event, relaxation occurs, and the halo overdensity mass behaves as a damped oscillator as a function of time before settling on a new equilibrium within several Gyr. We then measure the new overdensity mass and compare it to the mass just before ejection. Finally, we repeat the simulation a number of times with different values of $f_{\mathrm{ej}}$.
@@
Finally, some dependence on the halo profile is expected, as its shape should affect the outcome of relaxation. Repeating this experiment for haloes with different masses and concentrations (not shown), we find only a weak dependence on concentration, where haloes with higher concentrations lose slightly less additional mass. This is in line, qualitatively, with the results of Elbers et al. (2025). The FLAMINGO results presented in this work, including those with different cosmologies, imply that any profile dependence (along with any other dependencies) of the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation must indeed be small.
@@ 4.2 Predictions based on observed baryon fractions
Similar to SP(k) (Salcido et al., 2023), the resummation model directly maps observed baryon fractions and their uncertainties to the suppression of the matter power spectrum, though without assuming a functional form for the $f_{\mathrm{b}}-M_{\mathrm{500c}}$ relation. While baryon fractions measured within any overdensity radius can be used with our model, fractions within $R_{\mathrm{500c}}$ ($f_{\mathrm{b,500c}}$) are currently most readily available. Lacking precise observations at larger radii, like $R_{\mathrm{200m}}$, we can use the results of §3.5 to predict $f_{\mathrm{ret,200m}}$ from $f_{\mathrm{b,500c}}$.
@@
The left-hand panel of Fig. 15 shows corrected baryon fractions within $R_{\mathrm{500c}}$ for several FLAMINGO simulations. Also shown are observed baryon fractions for two compilations: Akino et al. (2022), in red, and Dev et al. (2024), in purple. The Akino et al. (2022) data was derived for an X-ray-selected sample of 136 galaxy groups and clusters (HSC-XXL). The version of the data shown here was scaled to the DESY3 cosmology, and a $\pm 1\sigma$ confidence interval was derived using Monte Carlo simulation (Salcido & McCarthy, 2025, private communication). The Dev et al. (2024) data was compiled by combining observations across a wide range of wavelengths, and includes cold gas, hot gas (X-ray) and stars. These authors used eROSITA X-ray data that was not X-ray selected, which yield substantially lower gas fractions than pre-eROSITA X-ray-selected data. The halo masses for this data set were converted from $M_{\mathrm{200c}}$ to $M_{\mathrm{500c}}$ by undoing the same conversion factor employed by the authors, $M_{\mathrm{500c}}/M_{\mathrm{200c}}\approx 0.69$. Quantiles for these baryon fractions were provided by the authors. For both data sets, we assume the DESY3 cosmological baryon fraction to convert the baryon fractions to corrected baryon fractions. We see that the Dev et al. (2024) baryon fractions are lower than those of Akino et al. (2022), implying stronger feedback, as the authors themselves have noted. Indeed, above $M_{\mathrm{500c}}\approx 10^{12.5}\,\mathrm{M_{\odot}}$ only the most extreme feedback models in FLAMINGO are consistent with the baryon fractions of Dev et al. (2024) within $\approx 1\sigma$, and below this halo mass scale none are.
We use these observed baryon fractions as input to the resummation model, linearly extrapolating them in $(\log_{10}M_{\mathrm{500c}},f_{\mathrm{bc,500c}})$ where needed, although we note that the most important mass range for predicting the suppression signal is covered by both data sets (see §3.4). In addition, we use the results of §3.5 to predict $f_{\mathrm{ret,200m}}$ from $f_{\mathrm{b,500c}}$. Finally, we probe the uncertainty regions around the mean observed baryon fractions in two ways: once by assuming the uncertainties are fully correlated, meaning we use the upper or lower edge of the envelope around the data as input to the model; and once by assuming no correlations, by drawing random mean halo baryon fractions in each halo mass bin independently, interpreting the uncertainty regions of the Dev et al. (2024) data as $1\sigma$ Gaussian uncertainties as well. In the latter case, we take into account the asymmetry of the uncertainties by using the 84th and 16th percentiles to estimate $1\sigma$ uncertainties independently above and below the data. The results are shown in the right-hand panel of Fig. 15. Since we did not use small-scale information here, we have decreased the range in $k$ compared to previous figures. We also expanded the vertical range to fully show the strong suppression predicted by the Dev et al. (2024) baryon fractions.
[…]
@@
One may wonder whether the resummation model could be inverted to yield baryon fractions, given a suppression signal, similar to what Salcido & McCarthy (2025) did for SP(k). These could then be directly compared to observations. However, our model treats the baryon fraction of every halo mass bin as independent, which gives it great flexibility, but does mean that there is also a strong amount of degeneracy when trying to invert the model. Therefore, we would need to assume a functional form for $f_{\mathrm{b}}$ as a function of $M_{\mathrm{500c}}$; we find that a sigmoid function, specifically the hyperbolic tangent function, provides a great fit to the simulation results for $M_{\mathrm{500c}}\gtrsim 10^{12.5}\,\mathrm{M_{\odot}}$. Even in this case, there is still some degeneracy between halo mass bins, and between the baryon fractions of different overdensity regions. If we additionally assume the mapping between $f_{\mathrm{ret,500c}}$ and $f_{\mathrm{b,200m}}$ from §3.4, we can break most of the remaining degeneracy and recover the baryon fractions of several FLAMINGO simulations. However, when minimising the difference between the simulated suppression signal and the model-predicted one as a function of the $f_{\mathrm{b,500c}}$ sigmoid’s parameters, the high-mass baryon fractions are sometimes driven to unrealistically large values, as the $f_{\mathrm{ret,200m}}$ values saturate at high $f_{\mathrm{ret,500c}}$ (see the left-hand panel of Fig. 10). We leave improving the model inversion to future work.

### v2

@@ 3.4 Contribution strength of halo baryon fractions and uncertainty propagation
Examining Fig. 8, we can immediately see that the signal is dominated by halo masses around $M_{\mathrm{500c}}=10^{14}\,\mathrm{M_{\odot}}$, in line with the findings of e.g. vanDaalen2020, Debackere2020, Salcido2023 and vanLoonvanDaalen2024. The impact of other halo mass bins drops off quickly from there, with haloes in the mass range $M_{\mathrm{500c}}=10^{14.5-15}\,\mathrm{M_{\odot}}$ still contributing significantly on scales $k\,{<}\,{10}\,h\mathrm{\,Mpc^{-1}}$, while haloes with masses $M_{\mathrm{500c}}>10^{15}\,\mathrm{M_{\odot}}$ contribute negligibly on all scales. On the low-mass end, the large-scale contribution diminishes sharply below $M_{\mathrm{500c}}\approx 10^{13}\,\mathrm{M_{\odot}}$, while that on small scales ($k\,{\gtrsim}\,{5}\,h\mathrm{\,Mpc^{-1}}$) drops off much more slowly.
[…]
In both Fig. 8 and Fig. 9, the baryon fractions in haloes in the lowest two mass bins considered, $M_{\mathrm{500c}}=10^{11-11.5}\,\mathrm{M_{\odot}}$ and $10^{10.5-11}\,\mathrm{M_{\odot}}$, contribute negligibly on all scales shown. In the case of Fig. 9, the contribution of these lowest-mass haloes may be underestimated, as most of them are not well-resolved enough to form stars and therefore expel no gas, which they otherwise might have – though Fig. 8 shows that their contribution is small either way. Regardless, the inclusion of these mass scales in the model is still relevant to better measure the non-halo component.
@@ 3.5 Resummation based only on $\boldsymbol{\Delta=}\mathbf{500c}$
While increasingly accurate measurements of mean halo baryon fractions within $R_{\mathrm{500c}}$ are already observationally available for different $M_{\mathrm{500c}}$ (e.g. Akino2022) or even $M_{\mathrm{200c}}$ (e.g. Dev2024, though with some extrapolation), measuring these precisely for larger regions is more challenging. Promising techniques that derive large-scale gas fractions from measurements of the kSZ (e.g. Hadzhiyska2025a) or tSZ (e.g. Dalal2025) signal, or those using fast radio bursts to probe the intergalactic medium (e.g. Lemos2025), may change this in the near future. In the meantime, it is worth investigating how well the suppression signal can be reproduced using only baryon fractions measured within $R_{\mathrm{500c}}$.
One of the results of Debackere2020 was that the suppression signal is sensitive to the density profile of hot gas outside $R_{\mathrm{500c}}$. In the terms of the resummation model, the results therefore depend on the retained mass fractions on scales that are larger than the baryon fractions were measured on. We confirm this finding by showing the difference between two extreme assumptions for where the removed mass goes in Fig. 10. On the left-hand side, we show the result of assuming that all mass removed from within $R_{\mathrm{500c}}$ stays within $R_{\mathrm{200m}}$, by setting $f_{\mathrm{ret},i,\mathrm{200m}}=1$ for all halo masses. Clearly, this is a poor assumption, and the suppression of power on large scales is strongly underestimated as a result. On the right-hand side, we instead assume that all mass removed from within $R_{\mathrm{500c}}$ also makes it out of $R_{\mathrm{200m}}$ (where it is again assumed to be distributed similar to matter already outside DMO haloes), which means that $f_{\mathrm{ret},i,\mathrm{200m}}=1-(1-f_{\mathrm{ret},i,\mathrm{500c}})b_{i,\mathrm{500c}}f_{\mathrm{M},i,\mathrm{500c}}/(b_{i,\mathrm{200m}}f_{\mathrm{M},i,\mathrm{200m}})$. While this assumption is closer to the results of the simulations, how well it works is significantly dependent on halo mass and the strength of feedback: when either very little or a lot of mass is removed from within $R_{\mathrm{500c}}$, this assumption overestimates the retained mass fraction within $R_{\mathrm{200m}}$, i.e. more mass is actually removed from within $R_{\mathrm{200m}}$ (or less of it is accreted) than just that ejected from $R_{\mathrm{500c}}$.
@@
Figure 12: Similar to the bottom row of Fig. 7, and again including stellar fractions within $R_{\mathrm{2500c}}$, but now instead of including large-scale baryon fractions, we extrapolate the small-scale information. Using the correlation between the retained fractions within $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$ (see Fig. 11 and equation 17), we predict $f_{\mathrm{ret},i,\mathrm{200m}}$ given $f_{\mathrm{ret},i,\mathrm{500c}}$. This only slightly reduces the accuracy of the model for $k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$, and allows the resummation model to be applied given only mean halo baryon fractions within $R_{\mathrm{500c}}$.
The correlation between both retained mass fractions is particularly strong at high halo masses (large $f_{\mathrm{ret,500c}}$, see Fig. 3), where the retained mass of the larger region can generally be predicted to within percent accuracy from that of the smaller region. As we go to lower halo mass scales, however, the regions are increasingly decorrelated, and accurately predicting $f_{\mathrm{ret},i,\mathrm{200m}}$ (left) or $f_{\mathrm{ret},i,\mathrm{5xr500c}}$ (right) becomes more sensitive to the strength of feedback and its implementation (thermal or jet AGN feedback). Somewhat surprisingly, the correlation is slightly tighter for the larger region, $r<5\times R_{\mathrm{500c}}$ – likely because any dependence on the halo profile is removed, due to the larger radius being a fixed multiple of the smaller. This hints that we might obtain even better results if we replaced $R_{\mathrm{200m}}$ by a fixed multiple of $R_{\mathrm{500c}}$ in the resummation model, but we leave this for future work. We note that equation (17) also works well for predicting $f_{\mathrm{bc},\Delta}$ from $f_{\mathrm{bc,500c}}$, and translating these into retained mass fractions through equation (9) to predict $f_{\mathrm{ret},\Delta}$ results in a comparable scatter around the true values as directly using equation (17) for $f_{\mathrm{ret},\Delta}$.
@@ 4 Discussion
The improved resummation model presented in this work is able to reproduce the $z=0$ matter power suppression signal for a wide range of FLAMINGO simulations, without any free parameters. Based only on their halo spherical overdensity masses and mean halo baryon fractions, for example $R_{\mathrm{500c}}$ and $R_{\mathrm{200m}}$ or similar, the model yields percent-level accuracy up to $k\,{\approx}\,{10}\,h\mathrm{\,Mpc^{-1}}$, and even to $k\,{\gtrsim}\,{20}\,h\mathrm{\,Mpc^{-1}}$ if gas and stellar fractions are known inside $R_{\mathrm{2500c}}$ as well. Similar accuracy can be obtained at redshifts up to at least $z=1$ by adding a single parameter, although our results imply that this parameter may be independent of baryonic feedback and could therefore be a fixed function of redshift (see §3.6). We will explore this further in future work. Even without this parameter, the resummation model achieves $\lesssim 3\%$ accuracy up to at least $z=1$ for all $k\,{\lesssim}\,{10}\,h\mathrm{\,Mpc^{-1}}$.
@@ 4.1 The origin of the retained mass fraction relation
Central to the resummation model is the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation, which, for a given halo overdensity region and redshift, accurately returns the mass ratio between a halo in a hydrodynamic simulation or in the real Universe and its DMO equivalent, as a function of its halo baryon fraction (see Fig. 2, 3 and Fig. 4, as well as Appendix C). This relation is seemingly universal, at least within this family of simulations, showing no significant dependence on cosmology, feedback strength, or halo mass (at least for $M_{\mathrm{500c}}\gtrsim 10^{12}\,\mathrm{M_{\odot}}$). The existence of this relation, and its slope exceeding unity for all but the most massive haloes, implies that when mass is lost due to feedback or other baryonic processes, a predictable amount of cold dark matter mass is removed from overdense regions as well. As mentioned early in §3.1, this is expected based on purely gravitational processes: after mass is removed from the centre, the density enclosed in the ejection radius will decrease, and so overdensity radii will also decrease, which means mass between the new and original overdensity radii will be “lost”. Additionally, the halo potential will be lessened, reducing the binding energy of the particles, allowing some to stream out of the halo. Afterwards, a gravitational exchange of energy between halo particles will take place (halo relaxation), which will cause additional particles to move outward, or even become unbound, while for others the binding energy will increase. It is however not clear a priori whether this should lead to a precisely predictable amount of (relative) mass loss.
@@
To test whether this picture is accurate and how much freedom it allows, we ran a series of idealised DMO simulations using the swift simulation code (Schaller2024). Using ICICLE (Drakos2017), we generate initial conditions for a $M_{\mathrm{200m}}=10^{14}\,\mathrm{M_{\odot}}$ exponentially-truncated NFW halo with concentration $c\approx 4.5$ using 138 500138\,500 dark matter particles, such that $10^{5}$ of these reside within $R_{\mathrm{200m}}$. We note that the choice of halo mass is unimportant, as the mass can be rescaled without affecting the experiment. Similarly, the exponential truncation helps to define a stable “edge” to the halo, but does not influence the outcome. After letting the halo settle into a steady state, we instantaneously remove some fraction $f_{\mathrm{ej}}$ of its $M_{\mathrm{200m}}$ mass, crudely mimicking AGN feedback ejecting gas. We do so starting from the particle at the halo centre and moving radially outward until the desired fraction is reached, then remove these particles from the simulation. [^14] We also tried choosing some inner radius, and removing a random subset of the particles inside it, corresponding to the mass fraction to be “ejected”. The results are indistinguishable. We then continue the simulation. Immediately after this idealised ejection event, relaxation occurs, and the halo overdensity mass behaves as a damped oscillator as a function of time before settling on a new equilibrium within several Gyr. We then measure the new overdensity mass and compare it to the mass just before ejection. Finally, we repeat the simulation a number of times with different values of $f_{\mathrm{ej}}$.
@@
Finally, some dependence on the halo profile is expected, as its shape should affect the outcome of relaxation. Repeating this experiment for haloes with different masses and concentrations (not shown), we find only a weak dependence on concentration, where haloes with higher concentrations lose slightly less additional mass. This is in line, qualitatively, with the results of Elbers2025. The FLAMINGO results presented in this work, including those with different cosmologies, imply that any profile dependence (along with any other dependencies) of the $f_{\mathrm{ret}}-f_{\mathrm{bc}}$ relation must indeed be small.
@@ 4.2 Predictions based on observed baryon fractions
Similar to SP(k) (Salcido2023), the resummation model directly maps observed baryon fractions and their uncertainties to the suppression of the matter power spectrum, though without assuming a functional form for the $f_{\mathrm{b}}-M_{\mathrm{500c}}$ relation. While baryon fractions measured within any overdensity radius can be used with our model, fractions within $R_{\mathrm{500c}}$ ($f_{\mathrm{b,500c}}$) are currently most readily available. Lacking precise observations at larger radii, like $R_{\mathrm{200m}}$, we can use the results of §3.5 to predict $f_{\mathrm{ret,200m}}$ from $f_{\mathrm{b,500c}}$.
The left-hand panel of Fig. 16 shows corrected baryon fractions within $R_{\mathrm{500c}}$ for several FLAMINGO simulations. Also shown are observed baryon fractions for two compilations: Akino2022, in red, and Dev2024, in purple. The Akino2022 data was derived for an X-ray-selected sample of 136 galaxy groups and clusters (HSC-XXL). The version of the data shown here was scaled to the fiducial FLAMINGO cosmology (DESY3, see Schaye2023), and a $\pm 1\sigma$ confidence interval was derived using Monte Carlo simulation (Salcido2025, private communication). The Dev2024 data was compiled by combining observations across a wide range of wavelengths, and includes cold gas, hot gas (X-ray) and stars. These authors used eROSITA X-ray data that was not X-ray selected, which yield substantially lower gas fractions than pre-eROSITA X-ray-selected data. The halo masses for this data set were converted from $M_{\mathrm{200c}}$ to $M_{\mathrm{500c}}$ by undoing the same conversion factor employed by the authors, $M_{\mathrm{500c}}/M_{\mathrm{200c}}\approx 0.69$. Quantiles for these baryon fractions were provided by the authors. For both data sets, we assume the DESY3 cosmological baryon fraction to convert the baryon fractions to corrected baryon fractions. We see that the Dev2024 baryon fractions are lower than those of Akino2022, implying stronger feedback, as the authors themselves have noted. Indeed, above $M_{\mathrm{500c}}\approx 10^{12.5}\,\mathrm{M_{\odot}}$ only the most extreme feedback models in FLAMINGO are consistent with the baryon fractions of Dev2024 within $\approx 1\sigma$, and below this halo mass scale none are.
We use these observed baryon fractions as input to the resummation model, linearly extrapolating them in $(\log_{10}M_{\mathrm{500c}},f_{\mathrm{bc,500c}})$ where needed, although we note that the most important mass range for predicting the suppression signal is covered by both data sets (see §3.4). In addition, we use the results of §3.5 to predict $f_{\mathrm{ret,200m}}$ from $f_{\mathrm{b,500c}}$. Finally, we probe the uncertainty regions around the mean observed baryon fractions in two ways: once by assuming the uncertainties are fully correlated, meaning we use the upper or lower edge of the envelope around the data as input to the model; and once by assuming no correlations, by drawing random mean halo baryon fractions in each halo mass bin independently, interpreting the uncertainty regions of the Dev2024 data as $1\sigma$ Gaussian uncertainties as well. In the latter case, we take into account the asymmetry of the uncertainties by using the 84th and 16th percentiles to estimate $1\sigma$ uncertainties independently above and below the data. The results are shown in the right-hand panel of Fig. 16. Since we did not use small-scale information here, we have decreased the range in $k$ compared to previous figures. We also expanded the vertical range to fully show the strong suppression predicted by the Dev2024 baryon fractions.
@@
One may wonder whether the resummation model could be inverted to yield baryon fractions, given a suppression signal, similar to what Salcido2025 did for SP(k). These could then be directly compared to observations. However, our model treats the baryon fraction of every halo mass bin as independent, which gives it great flexibility, but does mean that there is also a strong amount of degeneracy when trying to invert the model. Therefore, we would need to assume a functional form for $f_{\mathrm{b}}$ as a function of $M_{\mathrm{500c}}$; we find that a sigmoid function, specifically the hyperbolic tangent function, provides a great fit to the simulation results for $M_{\mathrm{500c}}\gtrsim 10^{12.5}\,\mathrm{M_{\odot}}$. Even in this case, there is still some degeneracy between halo mass bins, and between the baryon fractions of different overdensity regions. If we additionally assume the mapping between $f_{\mathrm{ret,500c}}$ and $f_{\mathrm{b,200m}}$ from §3.4, we can break most of the remaining degeneracy and recover the baryon fractions of several FLAMINGO simulations. However, when minimising the difference between the simulated suppression signal and the model-predicted one as a function of the $f_{\mathrm{b,500c}}$ sigmoid’s parameters, the high-mass baryon fractions are sometimes driven to unrealistically large values, as the $f_{\mathrm{ret,200m}}$ values saturate at high $f_{\mathrm{ret,500c}}$ (see the left-hand panel of Fig. 11). We leave improving the model inversion to future work.